OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (–2, 4); Slope: 5 A. y + 4 = 5x – 10 B. y + 4 = 5(x + 2) C. y – 4 = 5(x + 2) D. y – 4 = 5(x – 2)

hartnn (hartnn):

USING SLOPE-POINT FORM \(y-y_1=m(x-x_1)\) where m= slope = 5 here and point (-2,4) gives x1=-2,y1 = 4 just plug in values!

OpenStudy (anonymous):

so A, B, C, or D ??

hartnn (hartnn):

you plug in values and tell me what you get. we don't give direct answers here.

OpenStudy (anonymous):

i dont know how to plug in values! im bad at this badddd!!!

hartnn (hartnn):

plug in means to just replace if you plug in value for m =5, you get \(y-y_1=5(x-x_1)\) you plug in for x1 and y1.

OpenStudy (anonymous):

UGHHHH

hartnn (hartnn):

just try once, if you make mistake, i am here to correct.

OpenStudy (anonymous):

I DIID

hartnn (hartnn):

what you did ? post here.